Calculate the max acceleration when given amplitude as 20cm and acceleration as 4y(where y is displacement)
1
Expert's answer
2019-01-04T11:00:11-0500

The displacement

y(t)=A cos⁡〖(ωt)=20 cos⁡〖(ωt) cm〗 〗

The acceleration

a(t)=y^'' (t)=-Aω^2 cos⁡(ωt)=-ω^2 y(t)

So

ω=2 rad/s

Finally, the max acceleration

a_max=Aω^2=20 cm×(2 rad/s)^2=80 cm/s^2